# AgMMU: A Comprehensive Agricultural Multimodal Understanding and Reasoning Benchmark

### v1

@@ Abstract
We curate a dataset AgMMU for evaluating and developing vision-language models (VLMs) to produce factually accurate answers for knowledge-intensive expert domains. Our AgMMU concentrates on one of the most socially beneficial domains, agriculture, which requires connecting detailed visual observation with precise knowledge to diagnose, e.g., pest identification, management instructions, etc. As a core uniqueness of our dataset, all facts, questions, and answers are extracted from 116,231 conversations between real-world users and authorized agricultural experts. After a three-step dataset curation pipeline with GPT-4o, LLaMA models, and human verification, AgMMU features an evaluation set of 5,460 multiple-choice questions (MCQs) and open-ended questions (OEQs). We also provide a development set that contains 205,399 pieces of agricultural knowledge information, including disease identification, symptoms descriptions, management instructions, insect and pest identification, and species identification. As a multimodal factual dataset, it reveals that existing VLMs face significant challenges with questions requiring both detailed perception and factual knowledge. Moreover, open-source VLMs still demonstrate a substantial performance gap compared to proprietary ones. To advance knowledge-intensive VLMs, we conduct fine-tuning experiments using our development set, which improves LLaVA-1.5 evaluation accuracy by up to 3.1%. We hope that AgMMU can serve both as an evaluation benchmark dedicated to agriculture and a development suite for incorporating knowledge-intensive expertise into general-purpose VLMs.
Figure 1: AgMMU is a multimodal knowledge-intensive dataset with expertise data from the agricultural domain, for which vision-language models (VLMs) must observe the details of images and provide precise and fact-based answers. Originally derived from real-world user-expert conversations, AgMMU consists of five knowledge types, including insect identification, species identification, disease categorization, management instruction, and symptom description questions. From the question-answering perspective, AgMMU features 5,460 multiple-choice questions (MCQs) like conventional vision-language benchmarks [53] and open-ended questions (OEQs) like SimpleQA [50]. We also carefully curate an agricultural knowledge base with 205,399 pieces of information for foundation model fine-tuning. We hope AgMMU can benefit both knowledge-intensive VLMs and the social good.
@@ 1 Introduction
Recent progress in large language models (LLMs) and vision language models (VLMs) has demonstrated remarkable capabilities in general knowledge reasoning and understanding, as exemplified by great performance on a variety of multimodal reasoning benchmarks [53, 36, 20, 50, 34, 28]. However, existing benchmarks often emphasize general-domain tasks and may not fully capture the performance limitations of these models in more specialized, knowledge-intensive domains. Generating factually accurate answers has recently attracted attention in language models, such as SimpleQA [49] proposed by OpenAI, but it still calls for studies on multimodal models in knowledge-intensive domains.
@@
Agriculture, a cornerstone application of scientific biological knowledge, presents a particularly challenging domain and has been extensively investigated in the vision and machine learning community since early years [16, 27, 38, 15, 18, 13]. Unlike common tasks in the general domain, agricultural problems often require both precise visual analysis (e.g., identifying specific diseases, pests, or plant conditions) and extensive domain knowledge (e.g., treatment protocols, growing conditions, and management practices). Visual-based data are exceptionally crucial for plant management, because many serious plant problems (e.g., pests, disease, nutrient stress, water stress) are recognized visually and often with advice from remote experts. The stakes are also high in this scenario: accurate and timely agricultural diagnostics can mean the difference between crop survival and failure, directly impacting food security and farmer livelihoods. Despite the critical nature of this domain, we lack comprehensive benchmarks to evaluate and improve the agricultural reasoning abilities of VLMs.
@@
The challenge of creating such benchmarks is multifaceted. First, biological and agricultural tasks, where computer vision is often applied, are notably labor-intensive. They require expert-labeled data specific to the target task. Meanwhile, agricultural expertise is scarce and specialized, making it difficult to curate high-quality evaluation data  [8]. Second, real-world agricultural problems are inherently multimodal and require both visual and textual understanding. Third, we do not have a clear protocol for what defines a representative distribution of realistic agricultural questions to be evaluated. These challenges have left a significant gap in our understanding of how well current AI systems can handle real-world agricultural problems. At the same time, there is strong current interest in LLM-based question answering services for agriculture, with agricultural chatbots being released by several large and small companies [11, 5, 39, 6, 9] and research projects [2].
In this work, we introduce the AgMMU (Agricultural Multimodal Understanding and Reasoning Benchmark), the first comprehensive benchmark designed to evaluate the capabilities of multimodal foundation models on agricultural knowledge and visual reasoning. Our benchmark leverages 116,231 real-world conversations between 2013-2024 on the Internet, which offers one-to-one expert answers from Cooperative Extension/University staff and volunteers within participating Land-Grant institutions [14]. This long-form real-world question-answering data source ensures that our benchmark captures authentic challenges faced by farmers and gardeners, including complex visual symptoms, varied environmental conditions, and nuanced management decisions. We design a multi-step pipeline to construct our dataset, which is composed of conversation categorization, knowledge extraction, and evaluation question generation procedures. We leverage large multimodal foundation models [10, 1] to process the raw information. The final dataset consists of a wide range of plant-related questions and factual knowledge, including disease and insect identification, recognition of plant species, management instruction, and symptom description tasks. An illustration of our benchmark is shown in Figure 1.
@@
Evaluations conducted on a diverse collection of vision language models have revealed that existing large-scale foundation models are lacking in knowledge-extensive scenarios, suffering mainly from not having adequate knowledge to use nuanced situational context to conclude the correct answer. (refer to Sec. 4 for more details). Moreover, with our large multimodal agricultural knowledge base for fine-tuning, open-sourced model is able to achieve significant performance gain, exceeding strong closed-source models in almost all question subdomains, which highlight the significance of our curated.
Our contributions include: 1) AgMMU: A carefully curated evaluation set of 5,460 multiple-choice questions (MCQs) and open-ended questions (OEQs, where no candidate choices are provided), each extracted from QA pairs with expert answers, designed to test both visual understanding and knowledge application in agricultural contexts, and 2) AgBase-200K: A comprehensive agricultural multimodal knowledge dataset of 205,399 expert-verified facts and knowledge pairs, extracted from non-overlapping conversations, suitable for improving VLM performance on agricultural tasks. 3) A systematic evaluation of the leading VLMs on our benchmark, revealing critical insights about their current capabilities, limitations in handling knowledge-intensive agricultural queries, and systematic error analysis. Our analysis reveals a general phenomenon of VLMs in producing factually correct answers, and our AgMMU provides a complete training and evaluation suite for investigating such problems.
@@ 2 Related Work
Multimodal Foundation Models and Benchmarks.   The field of multimodal AI has witnessed remarkable progress in recent years, marked by increasingly sophisticated foundation models and comprehensive evaluation benchmarks. CLIP [32] marks early breakthroughs in large-scale multimodal foundation models and demonstrates strong zero-shot visual recognition capabilities. The emergence of VLMs has led to a new generation of models like LLaVA [24] that combine visual perception with sophisticated reasoning capabilities. Other work [46, 35, 25] advances the field by incorporating more careful dataset curation, better vision-language alignment, and stronger visual backbones. More recent developments have pushed the boundaries of scale and capability. Models like GPT-4 [1], Claude 3 [40], and Gemini [41, 33] have demonstrated remarkable general-purpose visual reasoning abilities.
The evolution of multimodal models has been accompanied by increasingly sophisticated evaluation benchmarks [22]. Early benchmarks like VQAv2 [17] and GQA [20] focus on basic visual question-answering capabilities. ScienceQA [34] specifically targeted scientific reasoning with visual components, while Eyes-Wide-Shut [47] evaluates the model’s ability to avoid hallucination. Recent benchmarks have emphasized real-world applications and complex reasoning. RealWorldQA [52] tests the ability of the models to handle practical, everyday visual queries. The MMMU [53] benchmark represents a comprehensive effort to evaluate multimodal understanding across multiple domains and task types. Unlike these methods, AgMMU presents the first comprehensive and challenging vision-language benchmark targeting agriculture, a domain that faces severe data scarcity and lacks domain-experts for large-scale dataset curation.
Computer Vision in Agriculture and Biology.   Computer vision has been extensively applied in the biological and agricultural domains [48, 54, 31, 3, 55], with significant datasets and benchmarks driving progress in species identification and disease classification [37, 12, 26, 51]. The iNaturalist dataset [18] marked a milestone by providing millions of species observations from citizen scientists, enabling the creation of a dataset and a deep neural network model (InsectNet) for the identification of pests [7, 13]. This was followed by more specialized collections like BioScan-1M/5M [15] which focused on microscopic biological images, and TreeOfLife-10M [38], which expands the data in a comprehensive phylogenetic framework. These datasets have facilitated numerous advances in automated species identification, disease detection, and biological image analysis. However, they cannot support complex real-world agriculture and gardening question answering.
@@
TreeOfLife [38] CLS -
SimpleQA [50] OEQ
MMMU [53] MCQ
Our dataset MCQ+OEQ
Table 1: A comparison between AgMMU and some representative agricultural and general-domain datasets. Our proposed AgMMU is the first agricultural-domain multimodal question-answering datasets with evaluation and training sets. Moreover, our emphasis on expert domain and OEQs, instead of MCQs, uniquely challenges the factual knowledge precision of VLMs. (OEQ: open-ended question, MCQ: multi-choice question.)
@@ 3 The AgMMU Benchmark
Figure 2: Starting from raw user-expert conversations, we design a three-step data curation pipeline with carefully designed prompts and filtering. (1) Step 1. We employ LLaMA-70B to categorize the conversation and filter out the samples that do not fall under our selected agriculture sub-domains. (2) Step 2. A larger LLaMA-405B model extracts key pieces of agricultural knowledge from the long-form user-expert conversation. (3) Step 3. These facts either go to the evaluation set or knowledge base. For evaluation, the multiple-choice questions (MCQs) and open-ended questions (OEQs) are both formatted by GPT4o assisted with human quality verification.
AgMMU, short for “Agricultural Multimodal Understanding”, is a specially curated benchmark aimed at multimodal understanding in the agricultural domain. In addition to the importance of agriculture-related research, our AgMMU is also a novel benchmark for knowledge intensive multimodal understanding of general vision-language models, since agriculture-related questions generally require precise comprehension of image details (e.g., identification of pests, diseases, or weeds) and accurate memorization of facts (e.g., providing management suggestions).
To support studies on agricultural and knowledge-intensive VLMs, our AgMMU constructs a benchmark that contains both an agricultural knowledge base for training and an evaluation set with multiple-choice and open-ended questions. All data for AgMMU are collected from the Internet between 2013 and 2024, where users with agricultural questions communicate with agricultural experts from cooperative extension and university staff within land grant institutions [14]. These questions cover a wide range of plant knowledge, including weed management, pest control, general growth advice, generic plant identification, environmental stress management, and nutrient deficiency management. Examples are demonstrated in Figure 1.
Our AgMMU comprises two sets. (1) An evaluation set aiming to rigorously assess the abilities of existing models. It consists of 5,460 multiple-choice questions (MCQs) with curated choices, similar to the format of MMMU [53], and open-ended questions (OEQs) focusing on knowledge in short phrases, similar to SimpleQA [50]. (2) A large-scale knowledge base AgBase-200K supporting the training and development of models. It contains 205,399 pieces of facts extracted from 108,231 high-quality user-expert conversations. We also demonstrate that this knowledge base can be used to fine-tune existing VLMs.
Compared with related datasets in Table 1, our AgMMU has multi-facet advantages for developing VLMs. Regarding the biological and agricultural benchmarks of iNaturalist [18] and TreeOfLife-BioCLIP [38], AgMMU supports question-answering evaluation, which is crucial for a real-world VLM. As for general vision-language understanding benchmarks like MMMU [53], our dataset shares a similar emphasis as SimpleQA [50] on expert knowledge and factuality for VLMs. The large-scale training set AgBase-200K also distinguishes our dataset from common evaluation benchmarks in enabling expert-domain fine-tuning. As a combination of the above properties, AgMMU can uniquely benefit the development of both agricultural expert VLMs and the factual precision of general VLMs.
@@ 3.1 Dataset Curation
Curation Pipeline.   Each real-world conversation we collect contains a series of diverse facts. To convert a conversation into atomic QA pairs suitable for the evaluation of VLMs, we design a three-step pipeline as in Figure 2. (1) Categorization. We first tag user-expert QAs with a horticulture or agriculture domain and a primary plant organism of interest. (2) Knowledge Extraction. Based on the domain tag, we further decompose each expert-user conversation into its most important pieces of agricultural knowledge. (3) QA Generation. This step is for the evaluation set. We convert extracted agricultural facts into multiple-choice questions (MCQs) or open-ended questions (OEQs).
@@
Step 1: Categorization.   Real-world agricultural questions come in a great variety with an open-ended nature. To extract clear agricultural knowledge, we must first account for the inherent complexity of these questions. Therefore, our first step is to tag each question with an agriculture sub-domain. The seven categories we decide on are [[[[disease advice”, “weed management”, “pests control”, “growing advice”, “environmental stress”, “nutrient deficiency”, “generic identification”, and other]]]]. We chose these categories based on the main categories designed by gardening and agriculture sites [42, 45, 44, 29, 30]. In this step, we filter out conversations in the “growing advice” and “other” sub-domains, as these are typically not image-related questions. Notably, the type of knowledge in user questions and expert answers often depends significantly on the question’s sub-domain, which we use to guide the next step, where prompts are selected according to the specific agricultural sub-domain.
@@
Step 2: Knowledge Extraction.   In step 2, we extract agricultural knowledge available for training and evaluating VLMs. Across all questions, the user-expert conversation generally comprises a combination of identification (e.g., disease, or plant species) and management instructions. For example, in Figure 2, the user provides species identification information and symptom description, while the expert recognizes the diseases and offers management suggestions. To systematically extract knowledge from these conversations, we organize agricultural knowledge into categories of “species identification”, “disease identification”, “symptom/visual description”, “management instructions”, “insect/pest identification.”
@@
To increase the reliability of information extraction, we leverage the agriculture sub-domain predicted in the previous step and design different prompt templates for each sub-domain. Our prompt leverages the following techniques for information extraction at higher qualities. (1) Our pipeline provides manually annotated in-context examples, significantly improving knowledge extraction quality and avoiding hallucination. (2) Our prompt explicitly verifies whether the expert has provided facts relevant to the target category, enabling the dynamic inclusion of available facts. We discover that LLMs [10, 1] can reliably complete the information extraction step. More details of our prompts and observations are available in the supplementary materials.
@@
Step 3: QA Generation.   For rigorous evaluation, AgMMU adopts multiple-choice questions (MCQs) for all the knowledge types and open-ended questions (OEQs) for facts that are in the form of short phrases, inspired by MMMU [53] and SimpleQA [50], respectively. Note that only a small balanced set of knowledge is used for evaluation, while the rest goes to the knowledge base AgBase-200K. We use GPT-4o [1] to format agricultural facts extracted from previous steps into QA pairs.
@@
We begin by carefully selecting a subset of samples for evaluation through a rigorous filtering and balancing process. (1) Filtering is done by removing the conversations without management information or issue identification, as we empirically find that the absence of such information indicates expert uncertainty. We also filter out questions with uncertain terms, such as “unclear” in the extracted information. (2) We remove questions with management instruction terms that direct users to an external source rather than giving them an actual answer, typically containing the words “lab,” “arborist,” etc. (3) To balance between different agriculture sub-domains (step 1) and knowledge types (step 2), we prioritize an even distribution across the types of questions, while maximizing the diversity of categories, as shown in Figure 3. This process extracts 8000 facts with high-quality and balanced distribution for evaluation.

### v2

# AgMMU: A Comprehensive Agricultural Multimodal Understanding Benchmark

@@ Abstract
We present AgMMU, a challenging real-world benchmark for evaluating and advancing vision-language models (VLMs) in the knowledge-intensive domain of agriculture. Unlike prior datasets that rely on crowdsourced prompts, AgMMU is distilled from 116,231 authentic dialogues between everyday growers and USDA-authorized Cooperative Extension experts. Through a three-stage pipeline: automated knowledge extraction, QA generation, and human verification, we construct (i) AgMMU, an evaluation set of 746 multiple-choice questions (MCQs) and 746 open-ended questions (OEQs), and (ii) AgBase, a development corpus of 57,079 multimodal facts covering five high-stakes agricultural topics: insect identification, species identification, disease categorization, symptom description, and management instruction. AgMMU has three key advantages:
  • Authentic & Expert-Verified: All facts, images, and answers originate from real farmer and gardener inquiries answered by credentialed specialists, ensuring high-fidelity agricultural knowledge.
  • Complete Development Suite: AgMMU uniquely couples a dual-format evaluation benchmark (MCQ and OEQ) with AgBase, a large-scale training set, enabling both rigorous assessment and targeted improvement of VLMs.
  • Knowledge-intensive Challenge: Our tasks demand the synergy of nuanced visual perception and domain expertise, exposing fundamental limitations of current general-purpose models and charting a path toward robust, application-ready agricultural AI.
Benchmarking 12 leading VLMs reveals pronounced gaps in fine-grained perception and factual grounding. Open-sourced models trail after proprietary ones by a wide margin. Simple fine-tuning on AgBase boosts open-sourced model performance on challenging OEQs for up to 11.6% on average, narrowing this gap and also motivating future research to propose better strategies in knowledge extraction and distillation from AgBase. We hope AgMMU stimulates research on domain-specific knowledge integration and trustworthy decision support in agriculture AI development.
Code: https://github.com/AgMMU/AgMMU
Data: https://huggingface.co/datasets/AgMMU/AgMMU_v1
@@
Figure 1: AgMMU is a multimodal agricultural dataset that challenges vision-language models (VLMs) to observe the details of images and provide factually precise answers. Derived from real-world conversations between users and authorized experts by USDA-funded Cooperative Extension, AgMMU covers five major agricultural knowledge types (demonstrated in five columns of the figure). AgMMU features 746 multiple-choice questions (MCQs) like conventional vision-language benchmarks [53] and the same number of open-ended questions (OEQs) like SimpleQA [50], all validated by human annotators. We also curate an agricultural knowledge base with 57,079 pieces of information for foundation model fine-tuning, extracted from experts’ answers. AgMMU can benefit both knowledge-intensive VLMs and the social good of agriculture.
@@ 1 Introduction
Recent progress in large language models (LLMs) and vision language models (VLMs) has demonstrated remarkable capabilities in general knowledge understanding, as exemplified by the outstanding performance on a variety of multimodal understanding benchmarks [53, 37, 21, 50, 35, 30]. However, existing benchmarks often emphasize general-domain tasks and may not fully capture the limitations of current models in more specialized, knowledge-intensive domains.
@@
Agriculture, a cornerstone application of scientific biological knowledge, presents a particularly challenging domain and has been extensively investigated in the vision and machine learning community since the early years [17, 29, 39, 16, 19, 14]. Unlike common tasks in the general domain, agricultural problems often require both precise visual analysis (e.g., identifying specific diseases, pests, or plant conditions) and extensive domain knowledge (e.g., treatment protocols, growing conditions, and management practices). Visual data is exceptionally crucial for agriculture, because many serious problems (e.g., pests, disease, nutrient stress, water stress) can only be recognized visually from nuanced appearances, such as the color and shape of leaves. The stakes are also high in this scenario: accurate and timely agricultural diagnostics can mean the difference between crop survival and failure, directly impacting food security and farmer livelihoods. Despite the critical nature of this domain, we lack comprehensive benchmarks to evaluate and improve the agricultural understanding abilities of VLMs, especially the concerns from real users.
@@
The challenge of creating such benchmarks is multifaceted. First, biological and agricultural tasks, where computer vision is often applied, are notably labor-intensive for data collection. They require expert-labeled data specific to the target task. Meanwhile, agricultural expertise is scarce and specialized, making it difficult to curate high-quality evaluation data [9]. Second, real-world agricultural problems are inherently multimodal requiring both visual and textual understanding, combining the visual observation with background information. Third, no clear protocol defines a representative distribution of realistic agricultural questions. These challenges have left a significant gap in our understanding of how well current AI systems can handle real-world agricultural problems [12, 6, 40, 7, 10, 2].
In this work, we introduce the AgMMU (Agricultural Multimodal Understanding Benchmark), the first real-world derived agricultural benchmark designed to evaluate the capabilities of multimodal foundation models. Our benchmark leverages 116,231 real-world conversations between 2013-2024 hosted by Cooperative Extension [15], which offers one-to-one conversations between real-world users capturing images from their own devices and professional agricultural experts funded by USDA to provide answers. This real-world data source ensures that our benchmark captures authentic challenges which farmers and gardeners face every day, including complex visual symptoms, varied environmental conditions, and nuanced management decisions.
For data curation, we design an automatic pipeline to extract the agricultural knowledge (as in Fig. 2) from long-form expert answers and then hire human annotators to verify the quality of the final question-answer pairs. Based on our observation of the real-world questions, the final dataset covers five major types of agricultural questions and knowledge: insect identification, species identification, disease categorization, management instruction, and symptom description, as in Fig. 1.
@@
Evaluations of a diverse collection of VLMs have revealed that existing foundation models struggle with knowledge-intensive agricultural scenarios, suffering mainly from insufficient knowledge. (refer to Sec. 4 for more details). Moreover, with our preparation of a large multimodal knowledge base, the open-sourced model can be notably improved with plain fine-tuning, even exceeding strong closed-source models in almost all question subdomains. Such observations signify the value of our agricultural benchmark and also encourage exploration of better strategies than simple fine-tuning.
Our contributions include:
  1. 1.
    AgMMU: A carefully curated evaluation set of 746 multiple-choice questions (MCQs) and the same number of open-ended questions (OEQs, where no candidate choices are provided) extracted from real-world conversations between users and USDA-funded experts and verified by our human annotators. Each extracted QA is designed to test both visual understanding and knowledge application in agricultural contexts.
  2. 2.
    AgBase: A comprehensive agricultural multimodal knowledge dataset of 57,387 pieces of facts extracted from the long-form experts answers, coming from non-overlapping conversations from the evaluation set. This is suitable for improving VLM performance on agricultural tasks.
  3. 3.
    A systematic evaluation of the leading VLMs reveals their limitations in handling knowledge-intensive agricultural queries, along with an error analysis. In the future, AgMMU will provide a complete training and evaluation suite for investigating such problems.
@@ 2 Related Work
Multimodal Foundation Models and Benchmarks.   The evolution of multimodal foundation models has been accompanied by increasingly sophisticated evaluation benchmarks [23] concentrating on a evolving set of problems significant for AI applications. Early benchmarks like VQAv2 [18] and GQA [21] focus on basic visual question-answering capabilities. ScienceQA [35] specifically targeted scientific reasoning with visual components, while Eyes-Wide-Shut [48] evaluates the model’s ability to avoid hallucination. Recent benchmarks have emphasized real-world applications and complex reasoning. RealWorldQA [52] tests the ability of the models to handle practical, everyday visual queries. The MMMU [53] benchmark represents a comprehensive effort to evaluate multimodal understanding across multiple domains and task types. On top of these benchmarks, AgMMU presents the first real-world oriented vision-language benchmark targeting agriculture, a domain that faces severe data scarcity and lacks domain experts for large-scale dataset curation.
AI in Agriculture.   AI has been extensively applied in the biological and agricultural domains [49, 55, 33, 3, 56, 54, 57], with significant datasets and benchmarks driving progress in species identification and disease classification [38, 13, 27, 51, 28], answering crop science questions [54, 36], and knowledge retrieval [5]. In terms of the multimodal perspective of agriculture, the iNaturalist dataset [19] marked a milestone by providing millions of species observations from citizen scientists, enabling the creation of a dataset and a deep neural network model (InsectNet) for the identification of pests [8, 14]. This was followed by more specialized collections like BioScan-1M/5M [16], which focused on microscopic biological images, and TreeOfLife-10M [39], which expands the data in a comprehensive phylogenetic framework. These datasets have facilitated numerous advances in automated species identification, disease detection, and biological image analysis. With the advance of vision-language models (VLMs) in recent years and farmers’ need for direct analysis of images, multimodal agricultural benchmarks emerge [54, 28] and evaluate the capabilities of VLMs to address agricultural problems. However, a major limitation of these benchmarks is that their images and questions are curated from agricultural experts instead of real-world users with agricultural difficulties, which cannot capture the distribution and complexities of real-world plant growing. From this perspective, our AgMMU fills the gap by building from the conversations between real-world users and experts.
@@
SimpleQA [50] OEQ
ScienceQA [35] MCQ
MMMU [53] MCQ
Our AgMMU MCQ+OEQ
(a) Comparison with general benchmarks.
Datasets Type Multimodal Real-world Factuality
@@
CROP [54] MCQ
CDDM [28] OEQ
Our AgMMU MCQ+OEQ
(b) Comparison with agricultural benchmarks.
Table 1: Objective of AgMMU. (a) AgMMU provides a multimodal benchmark in expert domains with a training set. With open-ended questions (OEQs), AgMMU emphasizes the factual accuracy of models, where a model has to recall precise facts without relying on options. (b) Besides, AgMMU is unique in leveraging questions from real-world users instead of recruited annotators. These properties make AgMMU a unique benchmark for advancing the VLMs for agriculture.
@@ 3 The AgMMU Benchmark
3.1 Overview
Objective.   We build AgMMU, short for “Agricultural Multimodal Understanding,” targeting the major challenges for VLMs to serve users in agricultural domains: how can VLMs provide precise factual knowledge for real-world questions? This objective leads to two critical design choices for our data curation. (1) Real-world distribution. The questions, images, and answers are derived from conversations between real users and experts, rather than being curated from the web or textbooks by annotators. (2) Factual questions. To evaluate the factual precision of models, we follow SimpleQA’s [50] style by requiring the VLMs to answer open-ended questions (OEQs) in short phrases, directly mentioning the key knowledge helpful for the users. In addition to open-ended questions, we provide multiple-choice questions (MCQs) to align with most multimodal benchmarks like MMMU [53]. Our OEQs and MCQs combination provides a comprehensive evaluation.
Data source.   We carefully select the data source to accomplish the above objectives. AgMMU is curated from 116,231 real-world conversations between 2013-2024 hosted by Cooperative Extension [15]. Such a data source offers the following critical advantages: (1) Real-world users: the questions and images are posted by real users who have agricultural questions; (2) Authorized experts: USDA funds the cooperative extension to authorize a group of professional agricultural experts from universities to answer these questions. Only certified experts are allowed to provide answers, which ensures the quality of data; (3) Eligibility for research: With the Cooperative Extension funded by USDA and being a non-profitable organization, the data can be open-sourced for research purposes under the CC-BY-SA4.0 License. Pre-processing is executed to remove personal and confidential information from our release dataset.
AgMMU Overview.   We design pipelines involving both automatic processing and human verification to turn raw conversations into high-quality QA pairs for agriculture knowledge evaluation. After filtering and processing, our AgMMU comprises of: (1) an evaluation set with 746 MCQ and the same number of OEQ questions covering the major agricultural knowledge types (as in Fig. 1); (2) a larger scale development set with 57,079 QA pairs formatted as open-ended QA, which is disjoint with the evaluation set and aimed to support the fine-tuning of VLMs for agricultural knowledge. We compare AgMMU with related AI and agricultural benchmarks in Table 1, marking its uniqueness.
@@ 3.2 Dataset Curation
Curation Pipeline.   Each real-world conversation is a multi-turn QA between the user and expert, containing one or several pieces of factual agricultural knowledge. We design a four-step pipeline to convert such a conversation into one QA pair suitable for the VLM evaluation as in Fig. 2. (1) Categorization. We first tag user-expert QAs with an agriculture domain and extract the primary plant organism of interest. (2) Knowledge Extraction. Based on the domain tag, we further decompose each expert-user conversation into its most important pieces of agricultural knowledge, where our prompts vary for different domains tagged in the previous step. (3) QA Generation. This step is for the evaluation set. We convert extracted agricultural facts into multiple-choice questions (MCQs) or open-ended questions (OEQs) for VLM evaluation. (4) Human Verification. We hire human annotators to filter out problematic QAs in the evaluation set.
@@
Step 1: Categorization.   Real-world agricultural questions come in a great variety with an open-ended nature. To extract clear agricultural knowledge, we must first account for the inherent complexity of these questions. Therefore, we first tag each question with an agriculture sub-domain. The seven categories we decide on are [[[[disease advice”, “weed management”, “pests control”, “growing advice”, “environmental stress”, “nutrient deficiency”, “generic identification", and other]]]], following the main categories designed by gardening and agriculture sites [43, 46, 45, 31, 32]. In this step, we filter out conversations in the “growing advice” and “other” sub-domains, as these are typically not image-related questions. Notably, the type of knowledge in user questions and expert answers often depends significantly on the question’s sub-domain, which we use to guide the next step, where prompts are selected according to the specific agricultural sub-domain.
Figure 2: Starting from raw user-expert conversations, we design a four-step data curation pipeline with carefully designed prompts and human verification. (1) We employ LLaMA-70B to categorize the conversation and filter out the samples that do not fall under our selected agriculture sub-domains. (2) A larger LLaMA-405B model extracts key agricultural knowledge from the long-form user-expert conversation. (3) These facts either go to the evaluation set or the development set. For evaluation questions, we utilize GPT4o to format the original QA and agricultural knowledge into multiple-choice questions (MCQs) and open-ended questions (OEQs). (4) Finally, human annotators verify the quality of the questions and only keep the qualified ones in the evaluation set.
@@
Step 2: Knowledge Extraction.   We extract agricultural knowledge for training and evaluating VLMs without introducing redundant languages from the original conversations, as in Fig. 2. Across all questions, the user-expert conversation generally comprises a combination of identification (e.g., disease, or plant species) and management instructions. For example, in Fig. 2, the user provides species identification information and symptom description, while the expert recognizes the diseases and offers management suggestions. To systematically extract knowledge, we organize the agricultural knowledge into categories of “species identification”, “disease identification”, “symptom/visual description”, “management instructions”, and “insect/pest identification.”
@@
To increase the reliability of information extraction, we leverage the agriculture sub-domain predicted in the previous step and design different prompt templates for each sub-domain. Our prompt leverages the following techniques: (1) Manually annotated in-context examples, which significantly improve knowledge extraction quality and avoiding hallucination. (2) Explicit fact verification, which determines whether the expert has provided relevant facts for the target category. We discover that LLMs [11, 1] can reliably complete the information extraction step. More details of our prompts and observations are available in the supplementary materials.
@@
Step 3: QA Generation.   For rigorous evaluation, AgMMU adopts multiple-choice questions (MCQs) and open-ended questions (OEQs), combining the advantages of MMMU [53] and SimpleQA [50], respectively. For the knowledge extracted from the previous step, a small and balanced subset is selected for evaluation, while the rest goes directly to AgBase for model development and fine-tuning. We use GPT-4o [1] to format agricultural facts into QA pairs.
@@
The following criteria guide the selection of the evaluation set via our close observation of the conversations. (1) Filtering is done by removing the conversations without management information or issue identification, as we empirically find that the absence of such information indicates expert uncertainty. We also filter out questions with uncertain terms, such as “unclear” in the extracted information. (2) We remove questions with management instruction terms that direct users to an external source rather than giving them an actual answer, typically containing the words “lab,” “arborist,” etc. (3) To balance between different agriculture sub-domains (step 1) and knowledge types (step 2), we prioritize an even distribution across the types of questions, while maximizing the diversity of categories, as shown in Fig. 3.
Then we employ GPT4o [1] to convert these facts into MCQs and OEQs, without losing the original intention of the user’s question. Depending on the knowledge types of the fact, we give the LLM a unique prompt to construct a corresponding question. We prompt it to generate three wrong answers and use the extracted fact as the ground truth to form one correct answer.
Step 4: Human verification   To guarantee the quality of the evaluation questions, we hire human annotators to manually verify the QA pairs from the following perspectives: (1) Faithfulness: whether the QA pair is faithful to the original conversation; (2) Certainty: whether the expert is certain about the answer; (3) Quality: are the images of good quality and correspond to the symptoms described by the user and the expert; (4) MCQ Feasibility: are all options other than the ground truth wrong. We strictly keep the data that pass all the evaluation criteria, resulting in 746 highest-quality MCQ and OEQ questions in AgMMU from 1742 samples.
Ethical Compliance.   We use automatic procedures assisted with human verification to erase the information that could leak the name, gender, address, or any other personal information for both users and experts. We will only distribute the processed AgMMU instead of the source data from the original conversation to protect the privacy of the users and experts. To the best of our verification, we have not found any human faces in the benchmark images.
3.3 Additional Properties of AgMMU